OpenStudy (anonymous):

can you satisfy me that how the triangular prism has 9 edges please and a cone has 1 edge?

OpenStudy (campbell_st):

well just draw then, or use google images... and search for them

OpenStudy (anonymous):



OpenStudy (campbell_st):

for the prism uses euler faces + edges = vertices + 2

OpenStudy (anonymous):

ok if edge means the way where two surfaces meet then in cone there is 1 edge and in triangluar prism how can we say that 5 faces or surfaces meeting to make 9 edges

ganeshie8 (ganeshie8):

edge is the line where two faces meet right ? and, each face has 4 sides

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

ofcourse triangle has 3 sides

OpenStudy (anonymous):

yeah edge means where 2 faces meet then in triangular prism faces are meeting to make 9 edges how become 9 edges

ganeshie8 (ganeshie8):

wats ur question

ganeshie8 (ganeshie8):

you're asking why 9 edges ?

OpenStudy (anonymous):

no i,m not asking why i'm asking how?

ganeshie8 (ganeshie8):

okay count the sides

ganeshie8 (ganeshie8):

u have two triangles and 3 rectangles # of sudes = 2x3 + 3x4 = 18

ganeshie8 (ganeshie8):

edge is the meeting point of faces, thus each edge eats up 2 sides :- number of edges = 18/2 = 9 if that looks plausible to u

OpenStudy (anonymous):

there are 9 sides but if we see a cone then?it has one edge

ganeshie8 (ganeshie8):

lets define terms :- side edge face vertex

ganeshie8 (ganeshie8):

side is a 2 dimensional term related to polygon all other 3 terms are polyhedra related terms ok

OpenStudy (anonymous):

oh sorry i know but edge is where two faces meet now i m doing what you say right away wait

ganeshie8 (ganeshie8):

take ur time, this is interesting topic to spend time in